# OminiControl: Minimal and Universal Control for Diffusion Transformer

### v4

@@ 3.3 Subjects200K datasets
Figure 6: Examples from our Subjects200Kdataset. Each pair of images shows the same object in varying positions, angles, and lighting conditions. The dataset includes diverse objects such as clothing, furniture, and animals, totaling over 200,000 images.
@@
Condition Model Method Controllability General Quality Text Consistency
[…]
Canny SD1.5 ControlNet 0.34 18.74 0.35 0.45 67.81 0.304
T2I-Adapter 0.22 20.06 0.35 0.39 67.88 0.305
FLUX.1 ControlNet 0.21 98.68 0.25 0.37 56.90 0.192
Ours 0.38 20.63 0.40 0.61 75.91 0.308
Depth SD1.5 ControlNet 923 23.02 0.34 0.47 70.73 0.308
T2I-Adapter 1560 24.72 0.27 0.39 69.99 0.309
FLUX.1 ControlNet 2958 62.20 0.26 0.38 66.84 0.212
Ours 903 27.26 0.39 0.55 75.06 0.308
Deblur FLUX.1 ControlNet 572 30.38 0.74 0.31 54.37 0.293
Ours 132 11.49 0.87 0.39 67.63 0.300
Colorization FLUX.1 ControlNet 351 16.27 0.64 0.43 70.95 0.278
Ours 24 10.23 0.73 0.43 70.74 0.301
Mask SD1.5 ControlNet 7588 13.14 0.40 0.41 67.22 0.300
FLUX.1 Ours 6248 15.66 0.48 0.45 72.61 0.306
Table 1: Quantitative comparison with baseline methods on five spatially aligned tasks. We evaluate methods based on Controllability (F1-Score for Canny, MSE for others), General Quality (FID, SSIM, MAN-IQA, MUSIQ), and Text Consistency (CLIP-Score). For F1-Score, higher is better; for MSE, lower is better. Best results are shown in bold.
To address this data challenge, we leverage a key observation: state-of-the-art DiT models like FLUX.1 [15] can generate identity-consistent image pairs when provided with appropriate prompts. Building on this, we develop an pipeline to create Subjects200K, a large-scale dataset specifically designed for subject-driven generation [^3] Detailed prompt design and synthesization pipeline are provided in the Appendix A.:
  • Prompt Generation. We use GPT-4o to generate over 30,000 diverse subject descriptions. Each description represents the same subject in multiple scenes. The goal is to highlight distinct objects or subjects while allowing natural variations.
  • Paired-image synthesis. We reorganized the collected descriptions into structured prompts. Each prompt describes the same subject in two different scenes. The template of such prompt is shown in Supplementary Material S3.These prompts are then fed into FLUX to generate image pairs. Each pair is designed to maintain subject consistency while varying in pose, lighting, and context.
@@
The resulting dataset (Figure 6) contains over 200,000 high-quality images spanning diverse categories including clothing, furniture, vehicles, and animals. Each subject appears in multiple contexts, providing rich training signals for learning robust subject-driven control. To facilitate future research, we release both the dataset and our complete generation pipeline [^4] Dataset and code available at: https://github.com/Yuanshi9815/Subjects200K..
@@ 4.1 Setup
Figure 7: Qualitative results comparing different methods. Left: Spatially aligned tasks across Canny, depth, out-painting, deblurring, colorization. Right: Subject-driven generation with beverage can, shoes and robot toy. Our method demonstrates superior controllability and visual quality across all tasks.
Implement details. Our method utilizes LoRA [4]for fine-tuning the base model with a default rank of 4. To preserve the model’s original capabilities and achieve flexibility, the LoRA scale is set to 0 when processing non-condition tokens by default.
@@
Baselines. For spatially aligned tasks, we compare our method with both the original ControlNet [42] and T2I-Adapter [22] on Stable Diffusion 1.5, as well as ControlNetPro [15], the FLUX.1 implementation of ControlNet. For subject-driven generation, we compare with IP-Adapter [41], evaluating its implementations FLUX.1 [39].
Evaluation metrics. We evaluate our model on both spatially aligned tasks and subject-driven generation. For spatially aligned tasks, we assess two aspects: generation quality and controllability. Generation quality is measured using FID [9], SSIM, MAN-IQA [40], and MUSIQ [11] for visual fidelity, along with CLIP Score [27] for semantic consistency. For controllability, we compute F1 Score between extracted and input edge maps in edge-conditioned generation, and MSE between extracted and original condition maps for other tasks (using Depth Anything for depth, color channel separation for colorization, etc.). For subject-driven generation, we propose a five-criteria framework evaluating both preservation of subject characteristics (identity preservation, material quality, color fidelity, natural appearance) and accuracy of requested modifications, with all assessments conducted through GPT-4o’s vision capabilities to ensure systematic evaluation. Detailed evaluation methodology is presented in Appendix B.1.
Evaluation protocol. We conducted evaluations on two datasets. For spatially aligned tasks, we use COCO 2017 validation set (5,000 images) resized to 512$\times$512, using task-specific conditions and associated captions as prompts with a fixed seed of 42. For subject-driven generation, we test on 750 text-condition pairs (30 subjects × 25 prompts) from DreamBooth[31] dataset with 5 different seeds, using one selected image per subject as the condition.
@@ 4.2 Main result
Spatially aligned tasks As shown in Table 1, we comprehensively evaluate our method against existing approaches on five spatially aligned tasks. Our method achieves the highest F1-Score of 0.38 on depth-to-image generation, significantly outperforming both SD1.5-based methods ControlNet [42] and T2I-Adapter [22], as well as FLUX.1-based ControlNetPro [15]. In terms of general quality metrics, our approach demonstrates consistent superiority across most tasks, showing notably better performance in SSIM [36], MAN-IQA [40], and MUSIQ [11] scores. For challenging tasks like deblurring and colorization, our method achieves substantial improvements: the MSE is reduced by 77% and 93% respectively compared to ControlNetPro, while the FID scores [9] improve from 30.38 to 11.49 for deblurring. The CLIP-Score metrics [27] indicate that our method maintains high text-image consistency across all tasks, suggesting effective preservation of semantic alignment while achieving better control and visual quality. As shown in Figure 7, our method produces sharper details and more faithful color reproduction in colorization tasks, while maintaining better structural fidelity in edge-guided generation and deblurring scenarios.
@@
Subject driven generation Figure 8 presents a comprehensive comparison against existing baselines. Our method demonstrates superior performance, particularly in identity preservation and modification accuracy. Averaging over random seeds, we achieve 75.8% modification accuracy compared to IP-Adapter (FLUX)’s 57.7%, while maintaining 50.6% identity preservation against IP-Adapter (SD 1.5)’s 29.4%. The advantage amplifies in best-seed scenarios, achieving 90.7% modification accuracy and 82.3% identity preservation - surpassing the strongest baselines by 15.8 and 18.0 percentage points, demonstrating effective subject-fidelity editing. These quantitative results are further corroborated by user studies presented in Appendix B.1.
@@
Comparative paramter efficiency. As shown in Table 2, our approach achieves remarkable parameter efficiency compared to existing methods. For the 12B parameter FLUX.1 model, our method requires only 14.5M trainable parameters (approximately 0.1%), which is significantly lower than ControlNet (27.5%) and IP-Adapter (7.6%). Even when utilizing the original VAE encoder from FLUX.1, our method still maintains high efficiency with just 0.4% additional parameters, demonstrating the effectiveness of our parameter-efficient design.
4.3 Empirical studies
Effect of training data.
Figure 9: Comparison of models trained with different data. The model trained by data augmentation tends to copy inputs directly, while model trained by our Subjects200Kgenerates novel views while preserving identity.
To validate the effectiveness of our Subjects200K dataset described in Section 3.3, we compare two training strategies for this task. The first approach relies on traditional data augmentation, where we apply random cropping, rotation, scaling, and adjustments to contrast, saturation, and color to the original images. The second approach utilizes our Subjects200K dataset. As shown in Figure 9, the model trained with data augmentation only learns to replicate the input conditions with minimal changes. In the first row, it simply places the taco plush toy in a bright room setting while maintaining its exact appearance and pose. Similarly, in the second row, the yellow alarm clock is reproduced with nearly identical details despite the window-side placement instruction. In contrast, our Subjects200K-trained model demonstrates the ability to generate diverse yet consistent views of the subjects while faithfully following the text prompts.
@@ 4.3 Empirical studies
Condition strength analysis. We evaluate our condition strength control through qualitative experiments. Figure 10 shows generated results with varying strength factor $\gamma\in{0.25,0.5,0.75,1.0}$. Results show that $\gamma$ effectively controls the generation process for both spatially aligned tasks like depth-to-image generation and spatially non-aligned tasks like subject-driven generation, enabling flexible control over the condition’s influence.
Figure 10: Comparison of models trained with different data. The model trained by data augmentation tends to copy inputs directly, while model trained by our Subjects200Kgenerates novel views while preserving identity.
@@
Impact of LoRA rank. We conducted extensive experiments with different LoRA ranks (1, 2, 4, 8, and 16) for the Canny-to-image task. As shown in Table 3, our experiments show that increasing the LoRA rank generally improves model performance, with rank 16 achieving the best results across multiple aspects: image quality (measured by FID and SSIM), condition control capability (measured by F1 Score), while maintaining competitive text-image consistency (measured by CLIP-Score). Notably, even with smaller ranks (e.g., rank 1), the model demonstrates competitive performance, especially in text-image alignment where it achieves the highest CLIP-Score of 0.765, showing the efficiency of our approach even with limited parameters.
@@
Conditioning depth. FLUX.1’s transformer architecture features two distinct types of blocks: early blocks that employ separate normalization modules for different modalities tokens (text and image) and later blocks that share unified normalization across all tokens. As shown in Table 3, experiments reveal that restricting condition signal integration to only these early blocks results in insufficient controllability over the generation process. This suggests that allowing the condition signals to influence the entire transformer stack is crucial for achieving the desired levels of control over the output. Notably, this finding indicates that the preview approaches[15, 39, 41, 42, 22] of inserting condition signals primarily in early blocks, which were effective in UNet-based architectures, may not fully translate to DiT-based models like FLUX.1.
@@ 5 Conclusion
OminiControl offers parameter-efficient image-conditioned control for Diffusion Transformers across diverse tasks using a unified token approach, without extra modules. Our method outperforms traditional approaches, and the new Subjects200K dataset—featuring over 200,000 high-quality, subject-consistent images—supports advancements in subject-consistent generation. Results confirm OminiControl’s scalability and effectiveness in diffusion models.
@@ Appendix B Additional experimental results
B.1 Evaluation for subject-driven generation
@@
B.2 Additional generation results

### v6

@@ 3.3 Subjects200K  datasets
To address this data challenge, we leverage a key observation: state-of-the-art DiT models like FLUX.1 [25] can generate identity-consistent image pairs when provided with appropriate prompts [20, 19]. Building on this, we develop an pipeline to create Subjects200K, a large-scale dataset specifically designed for subject-driven generation [^1] Detailed prompt design and synthesization pipeline are provided in the Section A of supplementary material.:
  • Prompt Generation. We use GPT-4o to generate over 30,000 diverse subject descriptions. Each description represents the same subject in multiple scenes.
  • Paired-image synthesis. We then reorganized the collected descriptions into structured prompts. Each prompt describes the same subject in two different scenes. The template of such prompt is shown in the Figure S3 of supplementary material.These prompts are then fed into FLUX to generate image pairs. Each pair is designed to maintain subject consistency while varying in context.
@@
The resulting dataset (Figure 6) contains over 200,000 high-quality images spanning diverse categories. Each subject appears in multiple contexts, providing rich training signals for learning robust subject-driven control. To facilitate future research, we release both the dataset and our complete generation pipeline [^2] Dataset and code available at supplementary material..
@@ 4.1 Setup
Figure 7: Qualitative comparison. Left: Spatially aligned tasks. Right: Subject-driven generation with beverage can and shoes. Our method demonstrates superior controllability and quality across all tasks. (More results are provided in supplementary materials.)
Implement details. Our method utilizes LoRA [8]for fine-tuning the base model with a default rank of 4. To preserve the model’s original capabilities and achieve flexibility, the LoRA scale is set to 0 by default when processing non-condition tokens.
@@
Baselines. For spatially aligned tasks, we compare our method with both the original ControlNet [61] and T2I-Adapter [37] on Stable Diffusion 1.5, as well as ControlNetPro [25], the FLUX.1 implementation of ControlNet. For subject-driven generation, we compare with IP-Adapter [60], evaluating its implementations FLUX.1 [58]. Additionally, we also with the official FLUX.1 Tools [24] implementation.
Evaluation of spatially aligned tasks. We evaluate our model on both spatially aligned tasks and subject-driven generation. For spatially aligned tasks, we assess two aspects: generation quality and controllability. Generation quality is measured using FID [16], SSIM, CLIP-IQA[53], MAN-IQA [59], MUSIQ [21], and PSNR[7] for visual fidelity, along with CLIP Text and CLIP Image scores [43] for consistency. For controllability, we compute F1 Score between extracted and input edge maps in edge-conditioned generation, and MSE between extracted and original condition maps for other tasks (using Depth Anything for depth, color channel separation for colorization, etc.). We use 5,000 images from COCO 2017 validation set, and resize them to 512$\times$512, then generate task-specific conditions and associated captions as prompts with a fixed seed of 42.
Evaluation of subject-driven generation. For subject-driven generation, we propose a five-criteria framework evaluating both preservation of subject characteristics (identity preservation, material quality, color fidelity, natural appearance) and accuracy of requested modifications, with all assessments conducted through GPT-4o’s vision capabilities to ensure systematic evaluation. Details are presented in the Section B.2 of supplementary material. We test on 750 text-condition pairs (30 subjects × 25 prompts) from DreamBooth[47] dataset with 5 different seeds, using one selected image per subject as the condition.
@@ 4.2 Main result
Spatially aligned tasks As shown in Table 2, we comprehensively evaluate our method against existing approaches on five spatially aligned tasks. Our method achieves the highest F1-Score of 0.38 on depth-to-image generation, significantly outperforming both SD1.5-based methods ControlNet [61] and T2I-Adapter [37], as well as FLUX.1-based ControlNetPro [25]. In terms of general quality metrics, our approach demonstrates consistent superiority across most tasks, showing notably better performance in SSIM [55], CLIP-IQA[53], MAN-IQA [59], MUSIQ [21] and PSNR[7] scores. For challenging tasks like deblurring and colorization, our method achieves substantial improvements: the MSE is reduced by 77% and 93% respectively compared to ControlNetPro, while the FID scores [16] improve from 30.38 to 11.49 for deblurring. The CLIP-Text and CLIP-Image metrics [43] indicate that our method maintains high consistency across all tasks, suggesting effective preservation of semantic alignment and image alignment while achieving better control and visual quality. As shown in Figure 7, our method produces sharper details and more faithful color reproduction in colorization tasks, while maintaining better structural fidelity in edge-guided generation and deblurring scenarios.
@@
Task Methods / Setting Base Model Controllability Image Quality Alignment
[…]
Canny ControlNet SD1.5 0.35 18.74 0.36 0.65 0.45 67.81 10.27 0.305 0.752
T2I Adapter 0.22 20.06 0.35 0.57 0.39 67.89 9.53 0.305 0.748
ControlNet Pro FLUX.1 0.21 98.69 0.25 0.48 0.37 56.91 9.22 0.192 0.537
Flux Tools 0.20 22.13 0.32 0.66 0.60 75.47 9.69 0.308 0.701
Ours 0.50 24.20 0.45 0.66 0.62 74.87 11.34 0.305 0.785
Depth ControlNet SD1.5 923 23.03 0.34 0.64 0.47 70.73 10.63 0.308 0.726
T2I Adapter 1560 24.72 0.28 0.61 0.39 69.99 9.50 0.309 0.721
ControlNet Pro FLUX.1 2958 62.20 0.26 0.55 0.39 66.85 9.38 0.212 0.547
Flux Tools 767 24.56 0.32 0.68 0.59 75.30 10.15 0.308 0.715
Ours 537 31.04 0.39 0.68 0.60 74.04 10.53 0.305 0.749
Mask ControlNet SD1.5 7588 13.14 0.68 0.58 0.42 67.22 18.96 0.300 0.848
Flux Tools FLUX.1 6610 11.40 0.73 0.56 0.45 68.92 18.37 0.305 0.874
Ours 6351 10.20 0.78 0.59 0.49 70.78 19.59 0.305 0.892
Colorization ControlNet Pro FLUX.1 994 30.38 0.75 0.40 0.31 54.38 16.23 0.279 0.781
Ours 73 10.37 0.92 0.56 0.48 69.40 21.56 0.305 0.884
Deblur ControlNet Pro FLUX.1 338 16.27 0.64 0.55 0.43 70.95 20.53 0.294 0.853
Ours 62 18.89 0.58 0.59 0.54 70.98 21.82 0.301 0.856
Table 2: Quantitative comparison with baseline methods on five spatially aligned tasks. We evaluate methods based on Controllability (F1-Score for Canny, MSE for others), Image Quality (SSIM, FID, CLIP-IQA, MAN-IQA, MUSIQ, PSNR), and Alignment (CLIP Text and CLIP Image). For F1-Score (used in Canny to Image task), higher is better; for MSE, lower is better. Best results are shown in bold.The second-best results are highlighted with underlines.
@@
Subject driven generation Figure 8 presents a comprehensive comparison against existing baselines. Our method demonstrates superior performance, particularly in identity preservation and modification accuracy. Averaging over random seeds, we achieve 75.8% modification accuracy compared to IP-Adapter (FLUX)’s 57.7%, while maintaining 50.6% identity preservation against IP-Adapter (SD 1.5)’s 29.4%. The advantage amplifies in best-seed scenarios, achieving 90.7% modification accuracy and 82.3% identity preservation - surpassing the strongest baselines by 15.8 and 18.0 percentage points, demonstrating effective subject-fidelity editing. These quantitative results are further corroborated by user studies presented in supplementary material B.2.
@@
Paramter efficiency As shown in Table 1, our approach achieves remarkable parameter efficiency compared to existing methods. For the 12B parameter FLUX.1 model, our method requires only 14.5M trainable parameters (approximately 0.1%), which is significantly lower than ControlNet (27.5%) and IP-Adapter (7.6%). Even when utilizing the original VAE encoder from FLUX.1, our method still maintains high efficiency with just 0.4% additional parameters, demonstrating the effectiveness of our parameter-efficient design.
@@
Condition strength factor We evaluate our condition strength control mechanism (Section 3.2.3) through qualitative experiments. Figure 9 shows generated results with varying strength factor $\gamma$. Results show that $\gamma$ effectively controls the generation process for both spatially aligned tasks like depth-to-image generation and non-aligned tasks like subject-driven generation, enabling flexible control over the condition’s influence.
Figure 9: Demonstration of the condition strength control.
4.3 Ablation Studies
To better understand the key factors that influence our model’s control capabilities, we conducted comprehensive ablation studies examining parameter efficiency, architectural decisions, and component contributions.
@@ 4.3 Ablation Studies
Impact of LoRA rank. We conducted extensive experiments with different LoRA ranks (1, 2, 4, 8, and 16) for the Canny-to-image task. As shown in Table 3, our experiments show that increasing the LoRA rank generally improves model performance, with rank 16 achieving the best results across multiple aspects. However, even with smaller ranks (e.g., rank 1), the model demonstrates competitive performance, especially in text-image alignment, showing the efficiency of our approach even with limited parameters.
@@
Conditioning depth. FLUX.1’s transformer architecture features two distinct types of blocks: early blocks that employ separate normalization modules for different modalities tokens (text and image) and later blocks that share unified normalization across all tokens. As shown in Table 3, experiments reveal that restricting condition signal integration to only these early blocks results in insufficient controllability. This suggests that allowing the condition signals to influence the entire transformer stack is crucial for generate better results. Notably, this finding indicates that the preview approaches[25, 58, 60, 61, 37] of injecting condition signals primarily in early blocks, which were effective in UNet-based architectures, may not fully translate to DiT-based models like FLUX.1.
Figure 10: Ablation study of critical modules for conditional generation. Given the censored Mona Lisa image and text prompt (bottom left), we test removing LoRA from different components.
Critical module analysis. To identify essential components for effective control, we conducted fine-grained ablation studies as shown in Figure 10. Results demonstrate that normalization layers and attention projections—particularly query ($W_{Q}$) and key ($W_{K}$)—are critical for maintaining control quality. Removing LoRA from these components significantly degrades conditional rendering, while value projections ($W_{V}$) have less impact. These insights reveal that control signals might propagate through normalization pathways and attention routing mechanisms rather than feature transformation processes.
@@ 5 Conclusion
OminiControl  offers parameter-efficient image-conditional control for DiT across diverse tasks using a unified token approach, requiring only 0.1% additional parameters. The Subjects200K dataset — featuring over 200,000 high-quality, subject-consistent images—further supports advancements in subject-driven generation, with experimental results confirming OminiControl ’s effectiveness in both spatially-aligned and non-aligned tasks.
However, the unified sequence approach increases the total number of tokens processed through the network, potentially limiting computational efficiency during inference. Addressing this token efficiency challenge while maintaining our method’s control capabilities represents an important direction for future research in parameter-efficient conditional generation.
@@ Appendix B Additional experimental results
B.1 Effect of training data
Figure S4: Comparison of models trained with different data. The model trained by data augmentation tends to copy inputs directly, while model trained by our Subjects200K generates novel views while preserving identity.
To validate the effectiveness of our Subjects200K  dataset described in Section 3.3, we compare two training strategies for this task. The first approach relies on traditional data augmentation, where we apply random cropping, rotation, scaling, and adjustments to contrast, saturation, and color to the original images. The second approach utilizes our Subjects200K  dataset. As shown in Figure S4, the model trained with data augmentation only learns to replicate the input conditions with minimal changes. In the first row, it simply places the taco plush toy in a bright room setting while maintaining its exact appearance and pose. Similarly, in the second row, the yellow alarm clock is reproduced with nearly identical details despite the window-side placement instruction. In contrast, our Subjects200K -trained model demonstrates the ability to generate diverse yet consistent views of the subjects while faithfully following the text prompts.
B.2 Evaluation for subject-driven generation
@@
B.3 Additional generation results